Using equations 1), 2) and 3), calculate the change in the enthalpy of the formation of methane, CH4, using chloroform, CHC13. CHC13 () + 3HCI (g) → CH4 (g) + 3C12 (g) 1) H2 (g) 2) Č (s) + 2H2 (g) HCl8) AH ° = -92.3 kJ CH4 (g) AH = -74.8 kJ + Cl2 (g) 3) C (s) + Hz1g) + Cl2) → CHC13 () 3) C (s) + H2 (g) + Cl2 (g) AH = -134.5 kJ
1
Expert's answer
2020-11-23T07:33:19-0500

1) ½ H2 (g) + ½ Cl2 (g) → HCl (g) ΔH° = -92.3 kJ


(2) C (s) + 2 H2 (g) → CH4 (g) ΔH° = -74.8 kJ


(3) C (s) + ½ H2 (g) + 3/2 Cl2 (g) → CHCl3(l) ΔH° = -134.5 kJ



In the formation of methane from chloroform, we need to reverse equation (3) since it has chloroform as a product, and we want it as a reactant. When you reverse this reaction, the sign of ΔH° will change from - to +.



(3) CHCl3 (l) → C (s) + ½ H2 (g) + 3/2 Cl2 (g) ΔH° = +134.5 kJ



Equation (2) will remain as is since methane is on the product side.



Reverse equation (1) since HCl has to be on the reactant side.


(1) HCl (g) → ½ H2 (g) + ½ Cl2 (g) ΔH° = +92.3 kJ


Multiply equation (1) by a factor of 3 since we need 3 moles of HCl in the formation of methane.


(1) [ HCl (g) → ½ H2 (g) + ½ Cl2 (g) ΔH° = +92.3 kJ ] x 3

So it becomes,

(1) 3 HCl (g) → 3/2 H2 (g) + 3/2 Cl2 (g) ΔH° = +276.9 kJ



We use all 3 equations, are


(1) 3 HCl (g) → 3/2 H2 (g) + 3/2 Cl2 (g) ΔH° = +276.9 kJ


(2) C (s) + 2 H2 (g) → CH4 (g) ΔH° = -74.8 kJ


(3) CHCl3 (l) → C (s) + ½ H2 (g) + 3/2 Cl2 (g) ΔH° = +134.5 kJ





3 HCl (g) + C (s) + 2 H2 (g) + CHCl3 (l) → 3/2 H2 (g) + 3/2 Cl2 (g) + CH4 (g) + C (s) + ½ H2 (g) + 3/2 Cl2 (g)



The net reaction after canceling and adding the coefficients of the common reactants / products will give:


CHCl3 (l) + 3 HCl (g) → CH4 (g) + 3 Cl2 (g)

ΔH° = +276.9 kJ - 74.8 kJ + 134.5 kJ

∆H° = +336.6 kJ



the change in the enthalpy of the formation of methane, CH4, using chloroform, CHC13, ∆H° = +336.6 kJ.